# Identifying and Analyzing Performance-Critical Tokens in Large Language Models

### v3

@@ Abstract
In-context learning (ICL) has emerged as an effective solution for few-shot learning with large language models (LLMs). However, how LLMs leverage demonstrations to specify a task and learn a corresponding computational function through ICL is underexplored. Drawing from the way humans learn from content-label mappings in demonstrations, we categorize the tokens in an ICL prompt into content, stopword, and template tokens. Our goal is to identify the types of tokens whose representations directly influence LLM’s performance, a property we refer to as being performance-critical. By ablating representations from the attention of the test example, we find that the representations of informative content tokens have less influence on performance compared to template and stopword tokens, which contrasts with the human attention to informative words. We give evidence that the representations of performance-critical tokens aggregate information from the content tokens. Moreover, we demonstrate experimentally that lexical meaning, repetition, and structural cues are the main distinguishing characteristics of these tokens. Our work sheds light on how large language models learn to perform tasks from demonstrations and deepens our understanding of the roles different types of tokens play in large language models.
Identifying and Analyzing Performance-Critical Tokens in
Large Language Models
Yu Bai1††thanks: Work in progress., Heyan Huang1, Cesare Spinoso-Di Piano2,3, Marc-Antoine Rondeau2, Sanxing Chen4, Yang Gao1 ††thanks: Corresponding author., and Jackie Chi Kit Cheung2,3,5 1Beijing Institute of Technology 2Mila – Quebec Artificial Intelligence Institute 3McGill University, 4Duke University 5Canada CIFAR AI Chair yubai@bit.edu.cn
@@ 1 Introduction
In-context learning (ICL) has become a popular technique employed with large language models (LLMs) (Brown et al., 2020). However, ICL has been shown to be unstable in that slight changes to the in-context prompts (e.g., reordering of demonstrations) can lead to substantial differences in performance (Lu et al., 2022; Zhang et al., 2022). This circumstance is difficult to control due to a lack of understanding of the model’s working mechanisms, leaving us uncertain about the exact process by which LLMs learn to infer a task specification from demonstrations and produce a computation function to implement that task specification. Previous papers have begun to explore this issue, focusing on specific aspects such as the label space (Min et al., 2022) and the hidden states of the last prompt token (Hendel et al., 2023; Todd et al., 2023), but have been limited in scope.
@@
In this work, we conduct a comprehensive study on how LLMs extract information that is valuable for improving task performance from demonstrations. Drawing from the way humans learn through content-label mappings in demonstrations, we categorize the tokens in an ICL prompt into content, stopword (Wilbur and Sirotkin, 1992), and template tokens, with the latter two typically ignored by humans due to their uninformative nature (Lenartowicz et al., 2014). We use content tokens as a category because they “are fixated (by human) about 85% of the time” (Rayner, 1998). With these categories in mind, we ablate the representations of different token types from the attention of ICL test examples, masking partial information during the model’s task-solving process, as shown in Figure 1. This ablation is intended to identify the types of tokens whose representations LLMs directly depend on to achieve high-level performance, thereby explaining how LLMs learn from demonstrations. We refer to these tokens that are critical for performance as performance-critical tokens.
@@
We show that template and stopword tokens (e.g., “Answer:”) are the most prone to be performance-critical tokens. In contrast, content tokens (e.g., “Union”, “player”, etc.) have a negligible impact on performance when their representations are eliminated from the attention of the test examples. This finding is counterintuitive since the original text of template and stopword tokens inherently do not possess any information found in the demonstrations. To explain this, we study the relationship among different types of tokens through ablation experiments that cut off the information flow between different kinds of tokens. We show that content tokens are indirectly leveraged by LLMs during ICL through aggregating their information into the representations of performance-critical tokens, raising questions about how LLMs use these tokens in general text processing and reasoning.
@@
Beyond identifying performance-critical tokens, we analyze them to better understand how they are leveraged by LLMs. Specifically, we investigate the characteristics which differentiate them from other tokens. We find the following three distinguishing characteristics: the lexical meaning of tokens as it relates to the task being solved, the repetition of tokens throughout the prompt, and the structural cues which the tokens provide to the prompt. Our findings indicate that the lexical meaning, repetition, and structural cues of performance-critical tokens contribute to task performance across all model sizes, suggesting that these characteristics contribute to the property of tokens being performance-critical.
@@
Our work reveals that we can identify and characterize the types of tokens whose representations are the most important in directly maintaining ICL task performance. Our results suggest that previous claims about ICL should be more nuanced, in that representations of tokens beyond label words (Wang et al., 2023) may also directly impact the task performance. We investigate the characteristics of lexical meaning, repetition, and structural cues related to performance-critical tokens which allow us to partially explain their importance and help us better understand how to avoid performance instability while using ICL. Additionally, our findings reinforce that we cannot assume LLMs and humans solve tasks in similar ways, emphasizing the need to develop better explanations of LLM behaviors. Overall, our research deepens the understanding of the roles different types of tokens play in LLMs, pointing to future work that explores how specific token representations can be leveraged for particular purposes (e.g., storing compressed information). Code and data will be released in the camera-ready version.
@@ 2 Related Work
2.1 Working mechanisms of ICL
Since the proposal of in-context learning (Brown et al., 2020), its working mechanisms have been extensively studied by the research community (Min et al., 2022; Liu et al., 2021; Olsson et al., 2022; Bhattamishra et al., 2023; Zhou et al., 2024). Min et al. (2022) suggest that demonstrations primarily provide the label space, the distribution of the input text, and the format of the sequence for the test example. They argue that the precise ground truth labels do not have significant importance. In contrast, Yoo et al. (2022) propose a differing view, stating that the impact of the ground truth labels depends on the experimental configuration. Mao et al. (2024) analyze in-context learning from the perspective of data generation. The perspective of supported training data is also leveraged to analyze ICL (Han et al., 2023). Zhao et al. (2024) propose to use coordinate systems to understand the working mechanism of in-context learning. Zhou et al. (2023) propose a comprehensive survey on the interpretation and analysis of in-context learning.
Our work investigates the working process of ICL in LLMs at inference time, demonstrating that certain specific tokens are more likely to possess representations that could affect the processing of final test samples, improving the task performance.
2.2 Function vectors of in-context learning
Todd et al. (2023) and Hendel et al. (2023) provide evidence that function vectors store information used to solve a task in ICL. They probe and extract the hidden representations of the final tokens in the prompt, which can then be added to or replace the corresponding vectors in a zero-shot example, yielding results comparable to using all demonstrations as context. Liu et al. (2023a) propose using an in-context vector to represent the target task and apply feature shifting to query examples. They feed each input and its target into an LLM, concatenate the latent states, and apply PCA to derive a vector aligned with the task. Finally, Wang et al. (2023) propose that label words in demonstrations act as information anchors, aggregating information from previous examples to provide it to the test example. This suggests label tokens may satisfy our definition of performance-critical tokens.
@@ 3.1 Notation
Component notation | Component example |
@@
Radio veteran Karmazin joins Sirius. Sirius Satellite Radio Inc. named former Viacom Inc. president Mel… | |
Business | |
Numbers point to NY. NEW YORK - The New York Yankees can achieve two milestones with one more victory… | |
Sports | |
ICL Prompt | Classify the news articles into the categories of World, Sports, Business, and Technology. |
Article: Radio veteran Karmazin joins Sirius. Sirius Satellite Radio Inc. named former Viacom Inc. president Mel… | |
Answer: Business | |
Article: Numbers point to NY. NEW YORK - The New York Yankees can achieve two milestones with one more victory… | |
@@
3.2 Definition of performance-critical token
In defining the performance-critical tokens, we measure the performance variation before and after incorporating the representations of specific tokens into the attention scope of the test example. Specifically, we define performance-critical tokens as the tokens that lead to both a noticeable performance improvement when their representations are included in the attention of test examples and performance degradation when they are excluded.
Formally, let be a large language model and be a classification dataset. We define as the set of representations of each token in the ICL prompt and as the set of representations of the test demonstration which is appended to for prediction (i.e., ). In addition, we let be some set of representations which may attend to from at inference time while performing ICL. For instance, would imply that, when is predicting the label of the test demonstration, the attention from the test example is restricted to the prompt’s instruction token representations.
@@ 4 Experimental settings
Datasets
we consider the most widely used text classification datasets used by previous studies (Zhao et al., 2021). For topic classification, we use the 4-way and 14-way datasets AGNews and DBPedia (Zhang et al., 2015). For textual entailment, we use the 3-way CB (De Marneffe et al., 2019) and 2-way RTE dataset (Dagan et al., 2005). We also use SST2 (Socher et al., 2013) and TREC (Voorhees and Tice, 2000) for sentiment and question classification tasks. Besides these classification tasks, we also present results in machine translation and question answering tasks to show that our findings can also be extended to text generation tasks. Results and analyses of these tasks are attached to Appendix L and Appendix M.
Evaluation
For each dataset, we randomly select 4 training demonstrations from the training set using 15 seeds, limited by the computational cost of LLM inference. For testing, we evaluate each setting on 500 randomly selected test examples. We show this sample size is sufficient by comparing results with 500 test examples and the full dataset using OpenLlama 3B and Llama 7B models, as shown in Appendix J. Instruction prompt is retained in all ablations, as it is essential for model performance (Yin et al., 2023). We use a fixed for all main results, with additional experiments using different provided in Appendix K, demonstrating that changing does not affect our main findings.
LLMs
We utilize the 7B, 13B, and 33B Llama models and a 3B OpenLlama model. Additional results with Llama 2 7B, Llama 2 13B, and Mistral 7B models are demonstrated in Appendix F, and models after supervised fine-tuning are tested in Appendix G. All experiments are conducted on a single A100 80G GPU. For the 13B and 33B models, we apply 8-bit quantization to fit them into a single GPU. The experiments are conducted using Huggingface Transformers (Wolf et al., 2020).
@@ 5.1 Token types
We categorize ICL tokens based on the structure of the ICL prompt, following our notation in Table 1. Firstly, we find it natural to categorize tokens based on the structure of ICL prompts where the tokens from the demonstration examples and the labels are separated by template tokens from and . Second, can be subdivided into content and stopword tokens, with the latter typically providing less useful information and often being ignored (Rayner, 1998) when humans use analogy to learn specific tasks. Guided by these intuitions, we categorize all the tokens in the ICL prompt into template tokens, stopword tokens, and content tokens as follows:
Template tokens (temp): In defining template tokens, we include all the tokens which serve as templates for the ICL prompt. This includes the tokens in and , as shown in Table 1.
Stopword tokens (stop): In defining stopword tokens, we include punctuation and conjunction words, such as [,], [.], etc., in the prompt. We use the stopword tokens appearing in the instructions111The stopword token list used in the main experiments and the ablation results with the complete NLTK (Loper and Bird, 2002) stopwords list are shown in Appendix H..
@@
Researchers might typically expect content tokens to be critical, as they account for most of the information in the demonstrations. However, in the following experiments, we find that the representations of template and stopword tokens have the greatest direct impact on performance.
@@ 5.2 Ablation on token types
To determine which token types are more likely to be performance-critical tokens whose representations directly affect the final performance significantly, we design two experiments which ablate representations or tokens based on token types. The first involves keeping and masking representations of different token types from the attention of the test example. The second involves dropping the various kinds of tokens from the ICL prompt. The main purpose of the first experiment is to identify the performance-critical tokens defined in Section 3.2, while the second experiment aims to cut off the information propagation of different types of tokens to further explore the functioning of performance-critical tokens. Illustrations of these two methods which we refer to as representation-level and token-level ablations are shown in Figure 1 and Figure 2. Examples for the representation-level ablation is provided in Appendix E.
Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
Open Llama 3B | Zero-shot | 22.0 | 20.0 | 23.6 | 5.4 | 44.4 | 1.8 | 19.5 |
cont | 26.2 | 52.1 | 30.1 | 7.4 | 51.9 | 37.9 | 14.8 | |
@@
temp | 56.5 | 86.7 | 27.1 | 62.2 | 56.4 | 52.3 | 37.4 | |
Llama 7B | Zero-shot | 25.0 | 29.2 | 41.4 | 0.0 | 54.2 | 3.6 | 25.6 |
cont | 32.4 | 57.9 | 42.5 | 12.5 | 55.5 | 46.1 | 15.6 | |
stop | 57.3 | 83.7 | 49.8 | 43.0 | 55.9 | 50.7 | 31.1 | |
temp | 70.8 | 90.2 | 58.4 | 66.2 | 66.3 | 73.5 | 45.3 | |
Llama 13B | Zero-shot | 59.0 | 18.0 | 37.0 | 0.0 | 0.0 | 0.0 | 19.0 |
cont | 27.7 | 52.4 | 33.5 | 10.9 | 61.7 | 41.7 | 19.0 | |
stop | 72.2 | 73.5 | 46.8 | 50.7 | 58.6 | 30.6 | 36.4 | |
temp | 80.0 | 92.3 | 58.6 | 76.9 | 68.5 | 47.7 | 51.7 | |
Llama 33B | Zero-shot | 70.2 | 88.6 | 60.6 | 30.2 | 58.1 | 19.6 | 54.6 |
cont | 24.4 | 61.7 | 62.1 | 10.5 | 65.2 | 63.6 | 6.7 | |
[…]
temp | 80.5 | 95.2 | 65.2 | 75.2 | 79.0 | 80.0 | 24.6 |
5.2.1 Representation-level ablation
Our first ablation stems from the intuition that if LLMs rely on the representations of specific token types to achieve high-level performance, they should perform adequately with only these representations. Performance should drop significantly if these tokens are removed from the test example’s attention. Hence, we first pass the entire ICL prompt to the LLM and then restrict the test example’s attention to representations of a particular token type (or types)222Since tokens have been shown to significantly impact performance (Wang et al., 2023), we always preserve the attention on the representations of the tokens. during its solving of the task. We compute task performances with every possible ablation combination, removing the representations of one (e.g., Standard ICL temp) or two token types (e.g., Zero-shot cont333Removing two types of tokens from Standard ICL is equivalent to adding the other type to Zero-shot.) from the attention of the test example. All the task performances and the averaged relative performance changes are reported, shown in Table 2 and 3. An illustration of this set of experiments is shown in Figure 1.
[…]
@@ 5.2.1 Representation-level ablation
Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
Open Llama 3B | Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 |
cont | 58.2 | 86.9 | 27.6 | 61.9 | 56.5 | 51.7 | 0.9 | |
stop | 51.8 | 78.9 | 28.8 | 30.3 | 53.6 | 45.2 | 9.9 | |
temp | 26.2 | 52.1 | 30.1 | 7.4 | 51.9 | 37.9 | 23.8 | |
Llama 7B | Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | 74.8 |
cont | 77.9 | 91.5 | 58.5 | 66.5 | 67.8 | 74.4 | 2.0 | |
stop | 78.5 | 88.7 | 39.3 | 66.7 | 60.6 | 60.4 | 9.1 | |
temp | 20.8 | 58.2 | 32.4 | 11.6 | 54.4 | 46.0 | 37.6 | |
Llama 13B | Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | 70.4 |
cont | 81.4 | 93.1 | 58.9 | 75.7 | 69.6 | 45.1 | 0.2 | |
stop | 79.8 | 85.8 | 64.4 | 73.6 | 64.5 | 40.2 | 2.4 | |
temp | 27.8 | 52.4 | 33.5 | 10.9 | 63.1 | 45.6 | 31.5 | |
Llama 33B | Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 |
cont | 82.3 | 95.4 | 64.9 | 76.1 | 80.4 | 82.0 | 1.5 | |
stop | 84.8 | 94.9 | 62.1 | 77.3 | 70.5 | 74.4 | 4.3 | |
temp | 24.4 | 61.7 | 60.6 | 10.5 | 67.7 | 68.5 | 32.7 |
5.2.2 Token-level ablation
In this section, we modify the ICL prompt by removing certain types of tokens from the ICL prompt555For template tokens, this includes both the tokens in the demonstrations and the test example to maintain their consistency. We included the analyses of only ablating the tokens in the demonstrations in Appendix P. to further investigate the relationship between different kinds of tokens, cutting off the information flow between the representations of different tokens, shown in Figure 2. When we ablate the template tokens, we preserve the answer and next-line tokens in the templates to maintain a basic separator between the demonstration inputs and outputs. Results averaged on all the datasets are presented in Figure 3. Detailed results on each dataset could be seen in Appendix O.
Our first finding from this ablation is that information is propagated from content token representations to performance-critical token representations, as shown by the contrast between representation-level and token-level ablation in Table 4. The representations of template and stopword tokens alone (i.e., Standard ICL cont in Figure 3) are less effective at affecting task performance, leading to worse performance than those where content token information is included in their attention (i.e., Standard ICL cont in Table 3).

### v4

@@ Abstract
In-context learning (ICL) has emerged as an effective solution for few-shot learning with large language models (LLMs). However, how LLMs leverage demonstrations to specify a task and learn a corresponding computational function through ICL is underexplored. Drawing from the way humans learn from content-label mappings in demonstrations, we categorize the tokens in an ICL prompt into content, stopword, and template tokens. Our goal is to identify the types of tokens whose representations directly influence LLM’s performance, a property we refer to as being performance-critical. By ablating representations from the attention of the test example, we find that the representations of informative content tokens have less influence on performance compared to template and stopword tokens, which contrasts with the human attention to informative words. We give evidence that the representations of performance-critical tokens aggregate information from the content tokens. Moreover, we demonstrate experimentally that lexical meaning, repetition, and structural cues are the main distinguishing characteristics of these tokens. Our work sheds light on how LLMs learn to perform tasks from demonstrations and deepens our understanding of the roles different types of tokens play in LLMs.
@@ 1 Introduction
In-context learning (ICL) has become a popular technique employed with large language models (LLMs) (brown2020language). However, ICL has been shown to be unstable in that slight changes to the in-context prompts (e.g., reordering of demonstrations) can lead to substantial differences in performance (lu-etal-2022-fantastically; zhang2022active). This circumstance is difficult to control due to a lack of understanding of the model’s working mechanisms, leaving us uncertain about the exact process by which LLMs learn to infer a task specification from demonstrations and produce a computation function to implement that task specification. Previous papers explored this issue, focusing on specific aspects such as the label space (min2022rethinking) and the hidden states of the last prompt token (hendel2023incontext; todd2023function), but have been limited in scope.
@@
In this work, we conduct a comprehensive study on how LLMs extract information that is valuable for improving task performance from demonstrations. Drawing from the way humans learn through content-label mappings in demonstrations, we categorize the tokens in an ICL prompt into content, stopword (wilbur1992automatic), and template tokens, with the latter two typically ignored by humans due to their uninformative nature (lenartowicz2014neurophysiological). We use content tokens as a category because they “are fixated (by human) about 85% of the time” (rayner1998eye). With these categories, we ablate the representations of different token types from the attention of ICL test examples, masking partial information during the model’s task-solving process, as shown in Figure 1. This ablation is intended to identify the types of tokens whose representations LLMs directly depend on to achieve high-level performance, thereby explaining how LLMs learn from demonstrations. We refer to these tokens that are critical for performance as performance-critical tokens.
@@
We show that template and stopword tokens (e.g., “Answer:”) are the most prone to be performance-critical tokens. In contrast, content tokens (e.g., “Union”, etc.) have a negligible impact on performance when their representations are eliminated from the attention of the test examples. This finding is counterintuitive since the original text of template and stopword tokens inherently do not possess any information found in the demonstrations. To explain this, we study the relationship among different types of tokens through ablation experiments that cut off the information flow between different kinds of tokens. We show that content tokens are indirectly leveraged by LLMs during ICL through aggregating their information into the representations of performance-critical tokens, raising questions about how LLMs use these tokens in general text processing and reasoning.
@@
Beyond identifying performance-critical tokens, we analyze them to better understand how they are leveraged by LLMs. Specifically, we investigate the characteristics which differentiate them from other tokens. We find the following three distinguishing characteristics: the lexical meaning of tokens as it relates to the task being solved, the repetition of tokens throughout the prompt, and the structural cues which the tokens provide to the prompt. Our findings indicate that the lexical meaning, repetition, and structural cues of performance-critical tokens contribute to task performance across all model sizes, suggesting that they contribute to the property of tokens being performance-critical.
@@
Our work reveals that we can identify and characterize the types of tokens whose representations are the most important in directly maintaining ICL task performance. Our results suggest that previous claims about ICL should be more nuanced, in that representations of tokens beyond label words (wang2023label) may also directly impact the task performance. We investigate the characteristics of lexical meaning, repetition, and structural cues related to performance-critical tokens which allow us to partially explain their importance and help us better understand how to avoid performance instability while using ICL. Additionally, our findings reinforce that we cannot assume LLMs and humans solve tasks in similar ways, emphasizing the need to develop better explanations of LLM behaviors. Overall, our research deepens the understanding of the roles different types of tokens play in LLMs, pointing to future work that explores how specific token representations can be leveraged for particular purposes (e.g., storing compressed information). Code and data are released at https://github.com/ybai-nlp/PCT˙ICL.
@@ 2 Related Work
Working mechanisms of ICL
Since the proposal of in-context learning (brown2020language), its working mechanisms have been extensively studied by the research community (min2022rethinking; liu2021makes; bhattamishra2023understanding; zhou2023mystery). min2022rethinking suggest that demonstrations primarily provide the label space, the distribution of the input text, and the format of the sequence for the test example. They argue that the precise ground truth labels do not have significant importance. In contrast, yoo2022groundtruth propose a differing view, stating that the impact of the ground truth labels depends on the experimental configuration. mao2024data analyze in-context learning from the perspective of data generation. The perspective of supported training data is also leveraged to analyze ICL (han2023understanding). zhao2024unveiling propose to use coordinate systems to understand the working mechanism of in-context learning. zhou2023mystery propose a comprehensive survey on the interpretation and analysis of in-context learning. lin2024dual model the dual modes of in-context learning (task retrieval vs. task learning) and phenomena like early performance drops. lilong introduce a long-context benchmark revealing degradation with very long demonstrations.
Our work investigates the working of ICL in LLMs at inference time, demonstrating that certain specific tokens are more likely to possess representations that could affect the processing of final test samples, improving the performance.
Function vectors of in-context learning
todd2023function and hendel2023incontext provide evidence that function vectors store information used to solve a task in ICL. They probe and extract the hidden representations of the final tokens in the prompt, which can then be added to or replace the corresponding vectors in a zero-shot example, yielding results comparable to using all demonstrations as context. liu2023incontext propose using an in-context vector to represent the target task and apply feature shifting to query examples. They feed each input and its target into an LLM, concatenate the latent states, and apply PCA to derive a vector aligned with the task. Finally, wang2023label propose that label words in demonstrations act as information anchors, aggregating information from previous examples. This suggests label tokens may satisfy our definition of performance-critical tokens. yang2025task show that compact task vectors naturally emerge and can be enhanced with an auxiliary loss. tikhonov2025one find that multiple vectors are often needed for complex tasks. dong2025understanding propose the Linear Combination Conjecture, explaining vector formation but noting limitations for high-rank tasks. jiangunlocking introduce function vectors to mitigate catastrophic forgetting in continual instruction tuning.
@@ 3 Preliminaries
Component example | ||
@@
| Radio veteran Karmazin joins Sirius. Sirius Satellite Radio Inc. named former Viacom Inc. president Mel… | |||
| Business | |||
| Numbers point to NY. NEW YORK - The New York Yankees can achieve two milestones with one more victory… | |||
| Sports | |||
Classify the news articles into the categories of World, Sports, Business, and Technology. | ||
| Article: Radio veteran Karmazin joins Sirius. Sirius Satellite Radio Inc. named former Viacom Inc. president Mel… | |||
| Answer: Business | |||
| Article: Numbers point to NY. NEW YORK - The New York Yankees can achieve two milestones with one more victory… | |||
@@
Definition of performance-critical token
In defining the performance-critical tokens, we measure the performance variation before and after incorporating the representations of specific tokens into the attention scope of the test example. Specifically, we define performance-critical tokens as the tokens that lead to both a noticeable performance improvement when their representations are included in the attention of test examples and a noticeable performance degradation when they are excluded.
Formally, let be a decoder-only transformer-based large language model (LLM) and be a classification dataset. We define as the set of representations of each token in the ICL prompt and as the set of representations of the test demonstration which is appended to for prediction (i.e., ). In addition, we let be some set of representations which may attend to from at inference time while performing ICL. For instance, would imply that, when is predicting the label of the test demonstration, the attention from the test example is restricted to the prompt’s instruction token representations.
@@ 4 Experimental settings
Datasets. We consider the most widely used text classification datasets used by previous studies (pmlr-v139-zhao21c). For topic classification, we use the 4-way and 14-way datasets AGNews and DBPedia (NIPS2015_250cf8b5). For textual entailment, we use the 3-way CB (de2019commitmentbank) and 2-way RTE dataset (dagan2005pascal). We also use SST2 (socher2013recursive) and TREC (voorhees2000building) for sentiment and question classification tasks. Besides these classification tasks, we also present results in machine translation and question answering tasks to show that our findings can also be extended to text generation tasks. Results and analyses of these generation tasks are attached to Appendix.
Evaluation. For each dataset, we randomly select 4 training demonstrations from the training set using 15 seeds, limited by the computational cost of LLM inference. For testing, we evaluate each setting on 500 randomly selected test examples. We show this sample size is sufficient by comparing results with 500 test examples and the full dataset using OpenLlama 3B and Llama 7B models, as shown in Appendix K. Instruction prompt is retained in all ablations, as it is essential for model performance (yin-etal-2023-read). We use a fixed for all main results of different models, with additional experiments with different provided in Appendix L, demonstrating that changing does not affect our main findings.
LLMs. We utilize the 7B, 13B, and 33B Llama models (touvron2023llama) and a 3B OpenLlama model. We also report additional results with Llama 2 7B, Llama 2 13B (touvron2023llama2), Mistral 7B (jiang2023mistral7b), and Gemma 3 4B (team2025gemma) models. Models after supervised fine-tuning are tested in Appendix H. All experiments are conducted on a single A100 80G GPU. For the 13B and 33B models, we apply 8-bit quantization to fit them into a single GPU.
@@ 5.1 Token types
We categorize ICL tokens based on the structure of the ICL prompt, following our notation in Table 1. Firstly, we find it natural to categorize tokens based on the structure of ICL prompts where the tokens from the demonstration examples and the labels are separated by template tokens from and . Second, can be subdivided into content and stopword tokens, with the latter typically providing less useful information and often being ignored (rayner1998eye) when humans use analogy to learn specific tasks. Guided by these intuitions, we categorize all the tokens in the ICL prompt into template tokens, stopword tokens, and content tokens as follows:
Template tokens (temp): In defining template tokens, we include all the tokens which serve as templates for the ICL prompt. This includes the tokens in and .
Stopword tokens (stop): In defining stopword tokens, we include punctuation and conjunction words, such as [,], [.], etc., in the prompt. We use the stopword tokens appearing in the instructions111The stopword token list used in the main experiments and the ablation results with the complete NLTK (loper2002nltk) stopwords list are shown in Appendix I..
@@
Researchers might typically expect content tokens to be critical, as they account for most of the information in demonstrations. However, in our experiments, we find that the representations of template and stopword tokens have the greatest direct impact on performance.
@@ 5.2 Ablation on token types
To determine which token types are more likely to be performance-critical tokens whose representations directly affect the final performance significantly, we design two experiments which ablate representations or tokens based on token types. The first involves keeping and masking representations of different token types from the attention of the test example. The second involves dropping the various kinds of tokens from the ICL prompt. The main purpose of the first experiment is to identify the performance-critical tokens defined in Section 3, while the second experiment aims to cut off the information propagation of different types of tokens to further explore the functioning of performance-critical tokens. Illustrations of these two methods which we refer to as representation-level and token-level ablations are shown in Figure 1 and Figure 2. Examples for the representation-level ablation is provided in Appendix F.
Representation-level ablation
Our first ablation stems from the intuition that if LLMs rely on the representations of specific token types to achieve high-level performance, they should perform adequately with only these representations. Performance should drop significantly if these tokens are removed from the test example’s attention. Hence, we first pass the entire ICL prompt to the LLM and then restrict the test example’s attention to representations of a particular token type (or types)222Since tokens have been shown to significantly impact performance (wang2023label), we always preserve the attention on the representations of the tokens. during its solving of the task. We compute task performances with every possible ablation combination, removing the representations of one (e.g., Standard ICL temp) or two token types (e.g., Zero-shot cont333Removing two token types from Standard ICL is equivalent to adding the other type to Zero-shot.) from the attention of the test example. All the task performances and the averaged relative performance changes are reported, shown in Table 2, 3, and 4. Results for Llama 7B and 13B models are shown in Appendix G. An illustration of this set of experiments is shown in Figure 1.
| Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. | ||
Zero-shot | 50.2 | 50.4 | 57.2 | 6.4 | 51.6 | 0.0 | 36.0 | ||
| cont | 0.9 | 61.0 | 50.6 | 12.9 | 48.7 | 53.2 | 1.9 | |||
| stop | 49.0 | 78.1 | 54.4 | 61.6 | 65.3 | 47.9 | 23.4 | |||
| temp | 81.1 | 82.6 | 55.2 | 65.5 | 63.9 | 55.4 | 31.3 | |||
Zero-shot | 56.2 | 90.8 | 49.0 | 7.6 | 70.0 | 46.4 | 53.3 | ||
| cont | 0.5 | 56.0 | 61.4 | 0.0 | 62.6 | 54.6 | 14.1 | |||
| stop | 47.2 | 76.8 | 65.2 | 65.3 | 66.5 | 56.8 | 9.6 | |||
| temp | 78.2 | 93.7 | 62.4 | 70.4 | 71.9 | 52.1 | 18.1 | |||
Zero-shot | 77.8 | 84.4 | 73.0 | 57.8 | 1.8 | 62.1 | 59.5 | ||
| cont | 43.3 | 52.0 | 66.6 | 10.1 | 64.3 | 60.7 | 10.0 | |||
| stop | 78.9 | 92.5 | 71.6 | 81.4 | 69.9 | 72.9 | 18.4 | |||
| temp | 81.7 | 95.9 | 63.9 | 83.3 | 77.9 | 72.5 | 19.7 | |||
Standard ICL | 85.0 | 93.2 | 58.3 | 66.7 | 66.3 | 55.0 | 70.7 | ||
| cont | 82.4 | 85.5 | 54.3 | 64.2 | 59.6 | 55.7 | 3.8 | |||
| stop | 84.8 | 88.0 | 51.7 | 65.7 | 65.8 | 53.2 | 2.5 | |||
| temp | 0.9 | 61.0 | 50.6 | 12.9 | 48.5 | 53.6 | 32.8 | |||
Standard ICL | 82.8 | 94.9 | 62.8 | 74.6 | 71.2 | 55.4 | 73.6 | ||
| cont | 79.0 | 94.1 | 62.7 | 72.4 | 72.1 | 53.6 | 1.3 | |||
| stop | 80.1 | 89.4 | 61.5 | 74.1 | 69.6 | 49.3 | 2.9 | |||
| temp | 0.5 | 56.0 | 61.4 | 0.0 | 68.2 | 54.3 | 33.5 | |||
Standard ICL | 82.2 | 97.0 | 67.4 | 82.4 | 73.6 | 78.6 | 80.2 | ||
| cont | 81.8 | 96.2 | 64.4 | 83.4 | 78.9 | 72.9 | 0.6 | |||
| stop | 81.3 | 97.0 | 66.5 | 80.5 | 75.5 | 75.7 | 0.8 | |||
| temp | 78.6 | 52.0 | 66.6 | 10.1 | 67.1 | 69.6 | 22.8 |
| Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. | ||
Zero-shot | 22.0 | 20.0 | 23.6 | 5.4 | 44.4 | 1.8 | 19.5 | ||
| cont | 26.2 | 52.1 | 30.1 | 7.4 | 51.9 | 37.9 | 14.8 | |||
| stop | 36.7 | 82.9 | 32.0∗ | 52.4 | 58.8 | 56.2 | 33.7 | |||
| temp | 56.5 | 86.7 | 27.1 | 62.2 | 56.4 | 52.3 | 37.4 | |||
Zero-shot | 70.2 | 88.6 | 60.6 | 30.2 | 58.1 | 19.6 | 54.6 | ||
| cont | 24.4 | 61.7 | 62.1 | 10.5 | 65.2 | 63.6 | 6.7 | |||
| stop | 72.9 | 92.7 | 66.7∗ | 69.1 | 69.6 | 63.0 | 17.7 | |||
| temp | 80.5 | 95.2 | 65.2 | 75.2 | 79.0 | 80.0 | 24.6 | |||
Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 | ||
| cont | 58.2 | 86.9 | 27.6 | 61.9 | 56.5 | 51.7 | 0.9 | |||
| stop | 51.8 | 78.9 | 28.8 | 30.3 | 53.6 | 45.2 | 9.9 | |||
| temp | 26.2 | 52.1 | 30.1 | 7.4 | 51.9 | 37.9 | 23.8 | |||
Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 | ||
| cont | 82.3 | 95.4 | 64.9 | 76.1 | 80.4 | 82.0 | 1.5 | |||
| stop | 84.8 | 94.9 | 62.1 | 77.3 | 70.5 | 74.4 | 4.3 | |||
| temp | 24.4 | 61.7 | 60.6 | 10.5 | 67.7 | 68.5 | 32.7 |
@@
| Setting | AGNews | SST2 | DBPedia | Avg. |
| Zero-shot | - | - | - | - |
| cont | 1.62 | 50.08 | 2.24 | 17.98 |
| stop | 77.32 | 87.22 | 76.72 | 80.42 |
| temp | 83.66 | 92.74 | 80.50 | 85.63 |
| Standard ICL | - | - | - | - |
| cont | 83.66 | 92.74 | 80.50 | 85.63 |
| stop | 84.22 | 94.14 | 79.50 | 85.95 |
| temp | 76.82 | 90.78 | 76.56 | 81.39 |
@@
Token-level ablation
In this section, we modify the ICL prompt by removing certain types of tokens from the ICL prompt555For template tokens, this includes both the tokens in the demonstrations and the test example to maintain their consistency. We included the analyses of only ablating the tokens in the demonstrations in Appendix Q. to further investigate the relationship between different kinds of tokens, cutting off the information flow between the representations of different tokens, shown in Figure 2. When we ablate the template tokens, we preserve the answer and next-line tokens in the templates to maintain a basic separator between the demonstration inputs and outputs. Results averaged on all the datasets are presented in Figure 3. Detailed results on each dataset could be seen in Appendix P.
Our first finding from this ablation is that information is propagated from content token representations to performance-critical token representations, as shown by the contrast between representation-level and token-level ablation in Table 5. The representations of template and stopword tokens alone (i.e., Standard ICL cont in Figure 3) are less effective at affecting task performance, leading to worse performance than those where content information is included in their attention (i.e., Standard ICL cont in Table 3).